OpenStudy (anonymous):

cos(x)-xsin(x)=0 solve for x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

you can only approximate

OpenStudy (anonymous):

or I think there is a limit

OpenStudy (anonymous):

rearrange cosx =xsinx 1/x = tan(x)

OpenStudy (anonymous):

now tan(x) -(1/x) =0

OpenStudy (anonymous):

ahh I dont know what to do from now, best you can do is approximate

OpenStudy (anonymous):

Thanks, this is the second derivative and I'm trying to find the inflection points. Wolframalpha gave me the answers, but not the steps needed.